# Taming Scalable Visual Tokenizer for Autoregressive Image Generation

### v1

@@ 4.1 Datasets and Metrics
The training of the visual tokenizer and autoregressive transformer are both on $256\times 256$ ImageNet [4]. For visual reconstruction, the reconstruction-FID, denoted as rFID [7], codebook utilization, and LPIPS [34] on ImageNet 50k validation set are adopted to measure the quality of reconstructed images. For visual generation, we measure the quality of image generation by the prevalent metrics FID, IS [21] and Precision/Recall [12].
@@ 4.2 Implementations Details
We adopt the same model architecture proposed in VQGAN [6]. The visual tokenizer is trained with the following settings: an initial $41e-$ 4 learning rate with $0.01$ multi-step decay mechanism, an Adam Optimizer [11] with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size with $330$ epochs, a combination of reconstruction, GAN [10], perceptual [34], commitment [6], entropy [32], double quantization losses, and LeCAM regularization [27] for training stability. When not specified, we use 16,384 codebook size, 256 code dimension, and 4 ResBlocks as our default tokenizer setting.
@@
It is anticipated that the plain autoregressive visual generation method can well exhibit the effectiveness of our visual tokenizer design. Specifically, we adopt Llama-based architecture where the RoPE [23], SwiGLU [22], and RMSNorm [35] techniques are incorporated. We further introduce AdaLN [18] for better visual synthesis quality. The class embedding serves as the start token and the conditional of AdaLN simultaneously. IBQ with width $w$, depth $d$ and head $h$ follows the scaling rules proposed in [24, 25], which is formulated as:
$w=64d,\quad h=d.$ (12)
All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer [15] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 450 training epochs corresponding the model size, gradient clipping of $1.0$, and $0.1$ dropout rate for input embedding, FFN module and conditional embedding.
@@ 4.3 Main Results
Tab. 1 shows the quantitative reconstruction comparison between IBQ and prevalent visual tokenizers. It can be seen that a significant drop in codebook usage on existing VQ methods when scaling codebook size (e.g., VQGAN [6] has a 44% usage for 1024 codebook size, while a 5.9% usage for 16,384 codebook size.), and code dimension (e.g., LlamaGen [24] has a 97% usage for 8-dimension codes, while a 0.29% usage for 256-dimension codes.) Therefore, the actual representational capacity is limited by the codebook collapse.
In contrast, our joint optimization of all codebook embedding and the visual encoder ensures the consistent distribution between them, facilitating the stable training of an large codebook size and embedding visual tokenizer with high utilization. Specifically, IBQ with 16,384 codebook size and 256 code dimension achieves 1.37 rFID, exceeding other VQ methods under the same downsampling rate and codebook size. By increasing the codebook size to 262,144, IBQ outperforms Open-MAGVIT2 [16] and achieves state-of-the-art reconstruction performance (1.00 rFID). We also perform a qualitative comparison with several representative VQ methods in Fig. 4. IBQ exhibits better visual soundness on complex scenarios such as faces and characters.
@@
In Tab. 7, we compare IBQ with other generative models, including Diffusion models, AR models, and variants of AR models (VAR [25] and MAR [14]) on class-conditional image generation task. With the powerful visual tokenizers of IBQ, our models show consistent improvements when scaling up the model size (from 300M to 2.1B), and outperform all previous vanilla autoregressive models at different scales of model size. Moreover, IBQ outperforms the diffusion-based model DiT [18], and achieves comparable results with the variants of AR models. These AR model variants focus on the architecture designs of transformers in the second stage, while our work is devoted to better visual tokenizers in the first stage. Therefore, we believe that with our stronger tokenizers, the AR models and their variants can be boosted further.
4.4 Scaling Up IBQ
[…]
IBQ shows promising scaling capacity in three aspects: 1) Codebook Size: As shown in Tab. 6, there is a significant improvement in reconstruction quality as the codebook size enlarges from 1024 to 16,384. Furthermore, IBQ enables high codebook utilization and consistent gain in visual soundness even training with 262,144 codes. 2) Model Size: Tab. 6 reveals that by extending the number of ResBlock both in the encoder and the decoder, the boost in reconstruction performance can be guaranteed. 3) Code Dimension: interestingly, we observe significant codebook usage increasements when scaling up code dimension. We assume that low-dimensional codes are less discriminative where similar ones tend to be clustered. It indicates that the representative codes are more likely to be selected under our global updating strategy. In contrast, codes with high-dimensional embedding are highly informative since they are mutually sparse in the representation space. Therefore, these codes can be evenly selected during training, which ensures high utilization with better performance. With the above factors, we realize a super large codebook of 262,144 codebook size and 256 dimensions with high codebook usage (84%), achieving the state-of-the-art reconstruction performance (1.00 rFID). To better illustrate the scaling properties, we also provide visualizations in Fig. 5.
@@ 4.5 Ablation Studies
For fair comparisons, we adopt LFQ [16] with 16,384 codes and replace its asymmetric token factorization with our vanilla transformer architecture. We compare with LFQ on both reconstruction and generation in Tab. 6, and our proposed IBQ performs better, which demonstrates increasing code dimension can improve the reconstruction ability of the visual tokenizer and further boost the visual generation.
@@ 5 Conclusion
In this paper, we identify the bottleneck in scaling tokenizers arising from the partial-update strategy in current VQ methods, which progressively broadens the distribution gap between encoded features and non-activated codes, ultimately leading to codebook collapse. To tackle this challenge, we propose a simple yet effective vector quantization method, coined as Index Backpropagation Quantization (IBQ), for scalable tokenizer training, which updates all codes by applying the straight-through estimator on the categorical distribution between visual features and all codebook embeddings, thereby maintaining consistent distribution between the entire codebook and encoded features. Experiments on ImageNet demonstrate that IBQ enables a high-utilization, large-scale visual tokenizer with improved performance in both reconstruction (1.00 rFID) and generation (2.05 gFID), confirming the scalability and effectiveness of our method.
@@
Appendix A Autoregressive Model Configurations
We show the detailed autoregressive model configurations in Tab. 8. We scale up the AR models from 342M to 2.1B following the scaling rules proposed in VAR [25].
Appendix B Comparison with Soft Vector Quantization
To comprehensively illustrate the rationality of our IBQ, we compare it with another global update method, Soft Vector Quantization dubbed as Soft VQ. During training, it adopts the weighted average of all code embeddings as the quantized feature $v_{q}$ and incorporates a cosine decay schedule of the temperature ranging from $0.9$ to $1e$-$6$ for one-hot vector approximation. As for inference, it switches back to the original VQGAN way, which selects the code with the highest probability for hard quantization.
Moreover, we provide an in-depth investigation by visualizing the distribution between the codebook and encoded features of Soft VQ. As shown in Fig. 6, although all-code updating strategy is enabled, the inappropriate quantization process tends to cluster codes mistakenly which subsequently suffers from low codebook usage (2.5%). We speculate that the force of the weighted average of code embeddings toward the encoded feature will smooth the codebook representation and result in similar and less informative code embeddings. In contrast, IBQ adopts hard quantization with index backpropagation. The hard quantization only involves the selected codes toward the encoded features for discriminative representation, thus ensuring precise quantization, while index backpropagation performs joint optimization of all codebook and the visual encoder to achieve consistent distribution. Considering the factors above, our proposed IBQ shows dominance in both reconstruction quality and codebook utilization.
Appendix C Additional Visualizations
Appendix D Limitation and Future Work
In this work, we mainly focus on visual tokenizers through scalable training, and adopt vanilla autoregressive transformers for visual generation. We believe that combining our powerful visual tokenizers with advanced AR models is a promising way to improve autoregressive visual generation. In addition, our models are trained on ImageNet [4], limiting the generalization ability. Pretraining tokenizers on more data and training AR models on text-image pairs would be helpful. We leave these for future work.
Model Parameters Width $w$ Head $h$ Depth $d$
IBQ-B 342M 16 16 1024
IBQ-L 649M 20 20 1280
IBQ-XL 1.1B 24 24 1536
IBQ-XXL 2.1B 30 30 1920
@@ Appendix D Limitation and Future Work
Figure 7: The upper part illustrates the IBQ tokenizer tested at 1024 $\times$ 1024 Unsplash. While the second part showcases the IBQ tokenizer tested at 256 $\times$ 256 Imagenet. (a) indicates the original images and (b) signifies the reconstructions.
@@
Figure 8: We showcases the 256 $\times$ 256 class conditional generation samples on Imagenet.

### v2

@@ 4.1 Datasets and Metrics
Both the visual tokenizers and autoregressive transformers are trained on $256\times 256$ ImageNet [9]. For reconstruction, we measure reconstruction-FID (rFID [13]), codebook utilization, and LPIPS [50] on the ImageNet 50k validation set. For generation, we evaluate image quality using generation FID (gFID), Inception Score [32], and Precision/Recall [19], following ADM evaluator [10].
@@ 4.2 Implementations Details
We adopt the same model architecture proposed in VQGAN [11]. The visual tokenizer is trained with the following settings: an initial $41e-$ 4 learning rate with $0.01$ multi-step decay mechanism, an Adam Optimizer [18] with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size with $330$ epochs, a combination of reconstruction, GAN [16], perceptual [50], commitment [11], entropy [47], double quantization losses, and LeCAM regularization [41] for training stability. Unless otherwise specified, we use a codebook size of 16,384, a code dimension of 256, and 4 ResBlocks as our default tokenizer setting.
@@
We use vanilla Autoregressive models ranging from 300M to 2.1B to validate the effectiveness of IBQ tokenizers in visual generation, adopting a Llama-based architecture with RoPE [36], SwiGLU [35], and RMSNorm [51]. AdaLN [28] is also incorporated for improved visual synthesis quality. The class embedding serves as both the start token and AdaLN condition. IBQ with width $w$, depth $d$ and head $h$ follows the scaling rules proposed in [37, 39], where $w=64d,h=d$. All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer [24] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 450 training epochs corresponding the model size, gradient clipping of $1.0$, and $0.1$ dropout rate for input embedding, FFN module and conditional embedding.
@@ 4.3 Main Results
Tab. 1 shows the quantitative reconstruction comparison between IBQ and prevalent visual tokenizers. Existing VQ methods show a significant drop in codebook usage when scaling codebook size (e.g., VQGAN [11] has a 44% usage for 1024 codebook size, while a 5.9% usage for 16,384 codebook size.), and code dimension (e.g., LlamaGen [37] has a 97% usage for 8-dimension codes, while a 0.29% usage for 256-dimension codes.) Therefore, the actual representational capacity is limited by the codebook collapse.
In contrast, IBQ’s joint optimization of codebook embeddings and the visual encoder maintains distribution consistency, enabling stable training of large-scale codebook with high utilization. Specifically, IBQ with 16,384 codebook size and 256 code dimension achieves 1.37 rFID, outperforming other VQ methods at the same downsampling rate and codebook size. Increasing the codebook size to 262,144, IBQ achieves state-of-the-art reconstruction with 1.00 rFID, surpassing Open-MAGVIT2 [25]. A qualitative comparison in Fig. 4 shows IBQ’s superior visual quality in complex scenarios such as faces and characters. Note that, we observe that incorporating additional facial data yields consistent improvements (see supplementary materials).
@@
In Tab. 7, we compare IBQ with other generative models, including Diffusion models, AR models, and variants of AR models (VAR [39] and MAR [23]) on class-conditional image generation. With powerful IBQ tokenizers, our models show consistent improvements when scaling up the model size (from 300M to 2.1B), and outperform all previous vanilla autoregressive models at different scales of model size. Moreover, IBQ outperforms the diffusion-based model DiT [28], and achieves comparable results with the variants of AR models. These AR model variants focus on the architecture designs of transformers in the second stage, while our work is devoted to better visual tokenizers in the first stage. Therefore, we believe that with our stronger tokenizers, the AR models and their variants can be boosted further.
4.4 Scaling Up IBQ Tokenizers
Existing VQ methods struggle to scale up due to the codebook collapse. For example, LlamaGen [37] sees a significant drop in usage and rFID when increasing the code dimension from 8 to 256 (97% $\rightarrow$ 0.29%, 2.19 rFID $\rightarrow$ 9.21 rFID), as shown in Tab. 1. This is due to their partial updates during training, which progressively widens the distribution gap between non-activated codes and encoded features.
Scaling Up Tokenizers Improves Reconstruction. IBQ tokenizers show promising scaling capacity for reconstruction in three aspects: 1) Codebook Size: As shown in Tab. 6, reconstruction quality improves significantly as the codebook size increases from 1,024 to 16,384, with high utilization and consistent visual soundness even at 262,144 codes. 2) Code Dimension: interestingly, we observe a notable increase in codebook usage when scaling code dimension in Tab. 6. We assume that low-dimensional codes are less discriminative and tend to be clustered, indicating that representative codes are more likely to be selected under our global updating strategy. In contrast, high-dimensional codes are highly informative due to their sparsity in the representation space, allowing for more even selection during training, which ensures high utilization with better performance. 3) Model Size: Tab. 6 reveals that by increasing the number of ResBlock both in both the encoder and decoder, a higher reconstruction performance can be guaranteed.
With these factors, we realize a super large codebook of 262,144 codebook size and 256 dimensions with high codebook usage (84%), achieving the state-of-the-art reconstruction performance (1.00 rFID). To better illustrate the scaling properties, we also provide visualizations in Fig. 5.
Figure 6: Scaling up IBQ tokenizers enables better generation, especially with larger autoregressive models (e.g., 1B param.).
Scaling Up Tokenizers Improves Generation. Scaling up IBQ tokenizers also enhances generation quality. As shown in Fig. 6, increasing the codebook size significantly improves reconstruction and generation FID, with a similar trend observed when scaling code dimensions. Moreover, with larger autoregressive models (e.g., 1B parameters), the improvement in generation quality becomes more remarkable, suggesting that scaling up generative models can further unlock the potential of IBQ tokenizers.
@@ 4.5 Ablation Studies
For fair comparisons, we adopt LFQ [25] with 16,384 codes and replace its asymmetric token factorization with our vanilla transformer architecture. Tab. 6 shows that IBQ outperforms LFQ in both reconstruction and generation, which demonstrates increasing code dimension can improve the reconstruction ability of the visual tokenizer and further boost the visual generation.
@@ 5 Conclusion
In this paper, we identify the bottleneck in scaling tokenizers (e.g., codebook size), stemming from the partial-update strategy in current VQ methods, which progressively enlarge the distribution gap between encoded features and non-activated codes, eventually leading to codebook collapse. To address this challenge, we propose a simple yet effective vector quantization method, termed as Index Backpropagation Quantization (IBQ), for scalable tokenizer training, which updates all codes by applying the straight-through estimator on the categorical distribution over visual features and all codebook embeddings, thereby maintaining consistent distribution between the entire codebook and encoded features. Experiments on ImageNet demonstrate that IBQ enables a high-utilization, large-scale visual tokenizer with improved performance in both reconstruction (1.00 rFID) and generation (2.05 gFID).
@@
Model Parameters Width $w$ Head $h$ Depth $d$ Lr Batch Size Epoch
IBQ-B 342M 16 16 1024 3e-4 768 300
IBQ-L 649M 20 20 1280 3e-4 768 350
IBQ-XL 1.1B 24 24 1536 3e-4 768 400
IBQ-XXL 2.1B 30 30 1920 3e-4 768 450
@@
Model Codebook Size Parameters Memory Time/epoch Usage
VQGAN 1,024 89.6M 19.5G 3h15min 44%
8,192 91.5M 19.7G 3h18min -
16,384 93.6M 19.8G 3h21min 5.3%
262,144 156M 21.2G 4h $\sim$0%
IBQ 1,024 89.6M 19.5G 3h20min 99%
8,192 91.5M 19.7G 3h30min 98%
16,384 93.6M 20G 3h40min 96%
262,144 156M 30.5G 9h 84%
Table 10: Training computational costs comparison between VQGAN and IBQ. (Tested on 8 A6000 gpus)
Appendix A Autoregressive Model Configurations
We show the detailed autoregressive model configurations and training settings in Tab. 8. We scale up the autoregressive models from 300M to 2.1B parameters, following the scaling rules proposed in VAR [39].
Appendix B Comparison with Soft Vector Quantization
To comprehensively illustrate the rationality of our IBQ, we compare it with another global update method, Soft Vector Quantization (Soft VQ). During training, it adopts the weighted average of all code embeddings as the quantized feature $v_{q}$ and incorporates a cosine decay schedule of the temperature ranging from $0.9$ to $1e$-$6$ for one-hot vector approximation. As for inference, it switches back to the original VQGAN way, which selects the code with the highest probability for hard quantization.
Moreover, we provide an in-depth investigation by visualizing the distribution between the codebook and encoded features of Soft VQ. As shown in Fig. 7, although all-code updating strategy is enabled, the inappropriate quantization process tends to cluster codes mistakenly, resulting in low codebook usage (2.5%). We speculate that the force of the weighted average of code embeddings toward the encoded feature will smooth the codebook representation and result in similar and less informative code embeddings. In contrast, IBQ adopts hard quantization with index backpropagation. The hard quantization only involves the selected codes toward the encoded features for discriminative representation, thus ensuring precise quantization, while index backpropagation performs joint optimization of the entire codebook and visual encoder to achieve consistent distribution. Considering the factors above, our proposed IBQ shows dominance in both reconstruction quality and codebook utilization.
Figure 8: Face reconstruction comparison. Scaling up tokenizers and finetuning tokenizers on face data can effectively improve facial reconstruction performance.
Method Ratio Codebook MS-COCO 2017 Imagenet-1k
[…]
LlamaGen 16 16384 8.40 20.28 0.55 2.47 20.65 0.54
Show-o 16 8192 9.26 20.90 0.59 3.50 21.34 0.59
Cosmos 16 64000 11.97 19.22 0.48 4.57 19.93 0.49
Open-MAGVIT2 16 16384 7.93 22.21 0.62 2.55 22.21 0.62
Open-MAGVIT2 16 262144 6.76 22.31 0.65 1.67 22.70 0.64
IBQ (Ours) 16 16384 7.67 21.58 0.62 2.06 22.01 0.61
IBQ (Ours) 16 262144 6.79 22.28 0.65 1.53 22.69 0.64
Table 11: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large-scale general-domain datasets and aim to serve text-conditional image generation. The results are reported under the same setup for fair comparison (text in gray signifies the results directly from Cosmos report). $\dagger$ indicates that LlamaGen loads the model initially trained on Imagenet while the others are training from scratch, i.e., MS-COCO and Imagenet-1k are excluded from training data.
Appendix C Training Costs
We evaluate the training costs of VQGAN and IBQ under varying codebook sizes using 8 A6000 GPUs. As shown in Tab. 10, the all-codes updating mechanism of IBQ incurs only a marginal increase in training costs compared to VQGAN when the codebook size is up to 16,384, yet it significantly improves codebook utilization. Specifically, IBQ introduces an additional 0.2 GB of memory usage and extends training time by 19 minutes, but increases codebook utilization from 5.3% to 96%. Furthermore, VQGAN fails to train with an extremely large codebook (i.e., 262,144 entries), whereas IBQ successfully achieves 84% utilization.
Appendix D Pretraining Tokenizer
We further unveil the representation capacity of our tokenizer by pretraining IBQ on large-scale domain datasets, i.e., 1) General: CapFusion [48], LAION-COCO [20], CC12M [8] and CC3M [34]. 2) High-quality: LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap, LAION-aesthetics [33], JourneyDB [27] and LAION-HD [^2] https://huggingface.co/datasets/yuvalkirstain/laion-hd-subset. We follow the same training settings stated in the manuscript while the training steps are $\sim$ 800,000. It can be seen in the Tab. 11 that IBQ achieves state-of-the-art performance compared to concurrent methods such as Cosmos [1], Show-o [45]. Although some recent efforts in residual tokenization [29, 12] can achieve better results, they are not listed here because residual techniques are orthogonal and compatible with IBQ. It is anticipated that our improvement on the naive quantization method better benefits the unified visual understanding and generation models compared to the residual one.
Appendix E Improving Face Reconstrution
Visual tokenizers trained on ImageNet may not perform as expected for face reconstruction. Increasing the codebook size can effectively mitigate this limitation. As shown in Fig. 8, increasing the codebook size from 16,384 to 262,144 leads to improved face reconstruction quality. Additionally, incorporating face data into the training set or fine-tuning on face-specific datasets are effective strategies for further enhancement. In particular, fine-tuning IBQ on the FFHQ dataset further enhances reconstruction performance.
Appendix F Additional Visualizations
@@
Figure 9: Reconstruction samples. The upper part illustrates the IBQ tokenizer tested at 1024 $\times$ 1024 Unsplash. While the second part showcases the IBQ tokenizer tested at 256 $\times$ 256 Imagenet. (a) indicates the original images and (b) signifies the reconstructions.
@@
Figure 10: Generation samples. We showcase the 256 $\times$ 256 class conditional generation samples on Imagenet.